# Few-Class Arena: A Benchmark for Efficient Selection of Vision Models and Dataset Difficulty Measurement

### v1

@@ Abstract
We propose Few-Class Arena (FCA), as a unified benchmark with focus on testing efficient image classification models for few classes. A wide variety of benchmark datasets with many classes (80-1000) have been created to assist Computer Vision architectural evolution. An increasing number of vision models are evaluated with these many-class datasets. However, real-world applications often involve substantially fewer classes of interest (2-10). This gap between many and few classes makes it difficult to predict performance of the few-class applications using models trained on the available many-class datasets. To date, little has been offered to evaluate models in this Few-Class Regime. We conduct a systematic evaluation of the ResNet family trained on ImageNet subsets from 2 to 1000 classes, and test a wide spectrum of Convolutional Neural Networks and Transformer architectures over ten datasets by using our newly proposed FCA tool. Furthermore, to aid an up-front assessment of dataset difficulty and a more efficient selection of models, we incorporate a difficulty measure as a function of class similarity. FCA offers a new tool for efficient machine learning in the Few-Class Regime, with goals ranging from a new efficient class similarity proposal, to lightweight model architecture design, to a new scaling law. FCA is user-friendly and can be easily extended to new models and datasets, facilitating future research work. Our benchmark is available at https://github.com/fewclassarena/fca.
@@ 1 Introduction
The de-facto benchmarks for evaluating efficient vision models are large scale with many classes (e.g. 1000 in ImageNet  [1], 80 in COCO  [2], etc.). Such benchmarks have expedited the advance of vision neural networks toward efficiency [3, 4, 5, 6, 7, 8, 9, 10] with the hope of reducing the financial and environmental cost of vision models [11, 12]. More efficient computation is facilitated by using quantization  [13, 14, 15], pruning  [16, 17, 18, 19], and data saliency [20]. Despite efficiency improvements such as these, many-class datasets are still the standard of model evaluation.
Real-world applications, however, typically comprise only a few number of classes (e.g, less than 10) [21, 22, 23] which we termed Few-Class Regime. To deploy a vision model pre-trained on large datasets in a specific environment, it requires the re-evaluation of published models or even retraining to find an optimal model in an expensive architectural search space [24].
@@
One major finding is that, apart from scaling down model and architectural design for efficiency, dataset difficulty also plays a vital role in model selection [25]  (described in Section 4.3).
Figure 1 summarizes several key findings under the Few-Class Regime. On the left graph in red are accuracy results for a range of number of classes $N_{CL}$ for what we call the “full model”, that is ResNet models pre-trained on the full 1000 classes of ImageNet (generally available from many websites). On the right are accuracy results for what we call “sub-models”, each of which is trained and tested on the same $N_{CL}$, where this number of classes is sampled from the full dataset down to the Few-Class Regime. Findings include the following. (a) Sub-models attain higher upper-bound accuracy than full models. (b) The range of accuracy widens for full models at few-classes, which increases the uncertainty of a practitioner selecting a model for few classes. In contrast, sub-models narrow the range. (c) Full models follow the scaling law [26]  in the dimension of model size - larger models (darker red) have higher accuracy from many to few classes. (d) Surprisingly, the scaling law is violated for sub-models in the Few-Class Regime (see the zoomed-in subplot) where larger models (darker blue) do not necessarily perform better than smaller ones (lighter blue). From these plots, our key insight is that, instead of using full models, researchers and practitioners in the Few-Class Regime should use sub-models for selection of more efficient models.
@@
We formally define Few-Class Regime as a scenario where the dataset has a limited number of classes. Real-world applications often comprise only a few number of classes (e.g. $N_{CL}<10$ or $10\%$ classes of a dataset). Consequently, Few-Class Arena refers to a benchmark to conduct research experiments to compare models in the Few-Class Regime. This paper focuses on the image classification task, although Few-Class Regime can generalize to object detection and other visual tasks.
@@
  • To be best of our knowledge, we are the first to explore the problems in the Few-Class Regime and develop a benchmark tool Few-Class Arena (FCA) to facilitate scientific research, analysis, and discovery for this range of classes.
@@
  • We incorporate dataset similarity as an inverse difficulty measurement in Few-Class Arena and propose a novel Silhouette-based similarity score named SimSS. By leveraging the visual feature extraction power of CLIP and DINOv2, we show that SimSS is highly correlated with ResNet performance in the Few-Class Regime with Pearson coefficient scores $\geq 0.88$.
@@ 2 Related Work
Visual Datasets and Benchmarks. To advance deep neural network research, a wealth of large-scale many-class datasets has been developed for benchmarking visual neural networks over a variety of tasks. Typical examples  [^1] A detailed list of many-class datasets used in this paper can be found in the Appendix. include 1000 classes in ImageNet [1] for image classification, and 80 object categories in COCO [2] for object detection. Previous benchmarks also extend vision to multimodal research such as image-text [27, 28, 29, 30]. While prior works often scale up the number of object categories for general purpose comparison, studies [31, 32] raise a concern on whether models trained on datasets with such a large number of classes (e.g. ImageNet) can be reliably transferred to real world applications often with far fewer classes. A close work to ours is vision backbone comparison [33] whose focus is on model architectures. Our perspective differs in a focus on cases with fewer number of classes, which often better aligns with real-world scenarios.
@@
Dataset Difficulty Measurement. Research has shown the existence of inherent dataset difficulty [32] for classification and other analytic tasks. Efficient measurement methods are proposed to characterize dataset difficulty using Silhouette Score [34], K-means Fréchet inception distance [35, 36, 37], and Probe nets [25]. Prior studies have proposed image quality metrics using statistical heuristics, including peak signal-to-noise ratio (PSNR) [38], structural similarity (SSIM) Index [39], and visual information fidelity VIF [40]. A neuroscience-based image difficulty metric [32] is defined as the minimum viewing time related to object solution time (OST) [41]. Another type of difficulty measure method consists of additional procedures such as c-score [42] , prediction depth [43] , and adversarial robustness [44] . Our work aligns with the line of research [45, 46, 47] involving similarity-based difficulty measurements: similar images are harder to distinguish from each other while dissimilar images are easier. Previous studies are mainly in the image retrieval context [48, 49, 50]. Similarity score is used in [51] with the limitation that a model serving similarity measurement has to be trained for one dataset. We push beyond this limit by leveraging large vision models that learn general visual features using CLIP [52] and DINOv2 [53]. The study [32] shows that CLIP generalizes well to both easy and hard images, making it a good candidate for measuring image difficulty. Supported by the evidence that better classifiers can act as better perceptual feature extractors [54] , in later sections we show how CLIP and DINOv2 will be used as our similarity base function.
@@
Despite the innovation of difficulty measure algorithms on many-class datasets, little attention has been paid to leveraging these methods in the Few-Class Regime. We show that, as the number of classes decreases, sub-dataset difficulty in the Few-Class Regime plays a more critical role in efficient model selection. To summarize, unlike previous work on many-class benchmarks and difficulty measurements, our work takes few-class and similarity-based dataset difficulty into consideration, and in doing so we believe the work pioneers the development of visual benchmark dedicated to research in the Few-Class Regime.
@@ 3 Few-Class Arena (FCA)
We introduce the Few-Class Arena (FCA) benchmark in this section. In practice, we have integrated FCA into the MMPreTrain framework [55], implemented in Python3 and Pytorch [^2] Code is available at https://github.com/fewclassarena/fca, including detailed documentation and long-term plans of maintenance..
3.1 Goals
@@ 3.3 Many-Class Full Dataset Trained Benchmark
Results are summarized in Table 1. We make several key observations: (1) models in different datasets (in rows) yield highly variable levels of performance by Top-1 accuracy; (2) no single best model (bold, in columns) exists across all datasets; and (3) model rankings vary across various datasets.
The first two observations are consistent with the findings in [25, 31]. For (1), it suggests there exists underlying dataset-specific difficulty. To capture this characteristic, we adopt the reference dataset classification difficulty number (DCN) [25] to refer to the empirically highest accuracy achieved in a dataset from a finite number of models shown in Table 1 and Figure 2 (a). For observation (3), we can examine the rankings among the ten models of ResNet50 and EfficientNet V2 in Figure 2 (b). ResNet50’s ranking varies dramatically for the different datasets, for instance ranking 7th on ImageNet1K and 1st on Quickdraw345. This ranking variability is also observed in other models (see all models in the Appendix). However, a common practice is to benchmark models – even for efficiency – on large datasets, especially ImageNet1K. The varied dataset rankings in our experiments expose the limitations of such a practice, further supporting our new benchmark paradigm, especially in the Few-Class Regime. In later sections, we leverage DCN and image similarity for further analysis.
@@
Table 1: Top-1 accuracy across ten models in ten datasets. Models are trained and tested on full datasets with their original number of classes (e.g. 1K from ImageNet1K); this is denoted in the last few digits of the abbreviation of the dataset name. The best score is highlighted in bold while the second best is underlined for each dataset. References for all models and datasets can be found in the Appendix.
@@
In the next subsections, we introduce three new types of benchmarks: (1) Few-Class, Full Dataset Trained Benchmark (FC-Full), which benchmarks vision models trained on the full dataset with the original number of classes; (2) Few-Class, Subset Trained Benchmark (FC-Sub), which benchmarks vision models trained on subsets of a fewer number of classes than the full dataset, and (3) Few-Class Similarity Benchmark (FC-Sim), which benchmarks image similarity methods and their correlation with model performance.
@@ 3.4 Few-Class Full Dataset Trained Benchmark (FC-Full)
Traditionally, a large number of models are trained and compared on many-class datasets. However, results for such benchmarks are not directly useful to the Few-Class Regime and many real-world scenarios. Therefore, we introduce the Few-Class Full Dataset Trained Benchmark (FC-Full), with the objective of effortlessly conducting large-scale experiments and analyses in the Few-Class Regime.
The procedure of FC-Full consists of two main stages. In the first stage, users select the models and datasets upon which they would like to conduct experiments. They can choose to download pre-trained model weights, which are usually available on popular model hubs (PyTorch Hub [56], TensorFlow Hub [57], Hugging Face [58], MMPreTrain [55] etc.). In case of no pre-trained weights available from public websites, users can resort to the option of training from scratch. To that end, our tool is designed and implemented to generate bash scripts for easily configurable and modifiable training through the use of configuration files.
In the second stage, users conduct benchmarking in the Few-Class Regime. By specifying the list of classes, Few-Class Arena automatically loads pre-trained weights of the chosen models and evaluates performance of the models on the selected datasets. Note that this process is accomplished through configuration files created by the user’s specifications, thus enabling hundreds of experiments to be launched by a single command. This dramatically reduces human effort that would otherwise be expended to run these experiments without Few-Class Arena.
@@ 3.6 Few-Class Similarity Benchmark (FC-Sim)
The intrinsic image difficulty of a dataset affects a model’s classification performance (and human) [59, 60, 32]. We show – as is intuitive – that the more similar two images are, the more difficult it is for a vision classifier to make a correct prediction. This suggests that the level of similarity of images in a dataset can be used as a proxy for a dataset difficulty measure. In this section, we first adopt and provide the basic formulation of similarity, the baseline of a similarity metric. Then we propose a Similarity-Based Silhouette Score to capture the characteristic of image similarity in a dataset.
We first adopt the basic similarity formulation from [51]. Intra-Class Similarity $S_{\alpha}^{(C)}$ is defined as a scalar describing the similarity of images within a class by taking the average of all the distinct class pairs in $C$, while Inter-Class Similarity denotes a scalar describing the similarity among images in two different classes ${C_{1}}$ and $C_{2}$. For a dataset $D$, these are defined as the mean of their similarity scores over all classes, respectively:
$S_{\alpha}^{(D)}=\frac{1}{|L|}\sum_{l\in L}S_{\alpha}^{(C_{l})}=\frac{1}{|L|\times|P^{(C_{l})}|}\sum_{l\in L}\sum_{i,j\in C_{l};i\neq j}\cos(\textbf{Z}_{i},\textbf{Z}_{j}),$ (1)
$S_{\beta}^{(D)}=\frac{1}{|P^{(D)}|}\sum_{a,b\in L;a\neq b}S_{\beta}^{(C_{a},C_{b})}=\frac{1}{|P^{(D)}|\times|P^{(C_{1},C_{2})}|}\sum_{a,b\in L;a\neq b}\sum_{i\in C_{1},j\in C_{2}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}).$ (2)
@@
To compensate the aforementioned drawback, we adopt the Silhouette Score (SS) [34, 61]: $SS(i)=\frac{b(i)-a(i)}{max({a(i),b(i)})}$, where $SS(i)$ is the Silhouette Score of the data point $i$, $a(i)$ is the average dissimilarity between $i$ and other instances in the same class, and $b(i)$ is the average dissimilarity between $i$ and other data points in the closest different class.
Observe that the above Intra-Class Similarity $S_{\alpha}^{(C)}$ already represents the tightness of the class $(C)$, therefore $a(i)$ can be replaced with the inverse of Intra-Class Similarity $a(i)=-S_{\alpha}(i)$. For the second term $b(i)$, we adopt the previously defined Inter-Class Similarity $S_{\beta}^{(C_{1},C_{2})}$ and introduce a new similarity score as Nearest Inter-Class Similarity $S^{\prime(C)}_{\beta}$, which is a scalar describing the similarity among instances between class $C$ and the closest class of each instance in $C$. The dataset-level Nearest Inter-Class Similarity $S^{\prime(D)}_{\beta}$ is expressed as:
$S^{\prime(D)}_{\beta}=\frac{1}{|L|}\sum_{l\in L}S^{\prime(C_{l},\hat{C_{l}})}_{\beta}=\frac{1}{|L|\times|P^{(C_{l},\hat{C_{l}})}|}\sum_{l\in L}\sum_{i\in C_{l},j\in\hat{C_{l}}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}).$ (3)
where $\hat{C}$ is the set of the nearest class to $C$ ($\hat{C}\neq C$). To summarize, we introduce our novel Similarity-Based Silhouette Score $SimSS$ [^4] The extended derivation is detailed in the Appendix.:

### v2

@@ Abstract
We propose Few-Class Arena (FCA), as a unified benchmark with focus on testing efficient image classification models for few classes. A wide variety of benchmark datasets with many classes (80-1000) have been created to assist Computer Vision architectural evolution. An increasing number of vision models are evaluated with these many-class datasets. However, real-world applications often involve substantially fewer classes of interest (2-10). This gap between many and few classes makes it difficult to predict performance of the few-class applications using models trained on the available many-class datasets. To date, little has been offered to evaluate models in this Few-Class Regime. We conduct a systematic evaluation of the ResNet family trained on ImageNet subsets from 2 to 1000 classes, and test a wide spectrum of Convolutional Neural Networks and Transformer architectures over ten datasets by using our newly proposed FCA tool. Furthermore, to aid an up-front assessment of dataset difficulty and a more efficient selection of models, we incorporate a difficulty measure as a function of class similarity. FCA offers a new tool for efficient machine learning in the Few-Class Regime, with goals ranging from a new efficient class similarity proposal, to lightweight model architecture design, to a new scaling law. FCA is user-friendly and can be easily extended to new models and datasets, facilitating future research work. Our benchmark is available at https://github.com/bryanbocao/fca.
@@ 1 Introduction
The de-facto benchmarks for evaluating efficient vision models are large scale with many classes (e.g. 1000 in ImageNet  (Deng et al., 2009), 80 in COCO  (Lin et al., 2014), etc.). Such benchmarks have expedited the advance of vision neural networks toward efficiency (Tan & Le, 2019a; Tan & L., 2021; Sinha & El-Sharkawy, 2019; Sandler et al., 2018; Howard et al., 2019; Iandola et al., 2016; Ma et al., 2018; Mehta & Rastegari, ) with the hope of reducing the financial and environmental cost of vision models (Patterson et al., 2021; Rae et al., 2021). More efficient computation is facilitated by using quantization  (Gysel et al., 2018; Han et al., 2015; Leng et al., 2018), pruning  (Cheng et al., 2017; Blalock et al., 2020; Li et al., 2016; Shen et al., 2022), and data saliency (Yeung et al., 2016). Despite efficiency improvements such as these, many-class datasets are still the standard of model evaluation.
Real-world applications, however, typically comprise only a few number of classes (e.g, less than 10) (Shao et al., 2020; A. Delplanque, 2022; Cai et al., 2021) which we termed Few-Class Regime. This introduces a crucial research question: what is the simplest baseline model capable of meeting performance criteria within this Few-Class Regime? To deploy a vision model pre-trained on large datasets in a specific environment, it requires the re-evaluation of published models or even retraining to find an optimal model in an expensive architectural search space (Scheidegger et al., 2019).
@@
One major finding is that, apart from scaling down model and architectural design for efficiency, dataset difficulty also plays a vital role in model selection (Scheidegger et al., 2021)  (described in Section 4.3).
Figure 1 summarizes several key findings under the Few-Class Regime. On the bottom left graph in red are accuracy results for a range of number of classes $N_{CL}$ for what we call the “full model”, that is ResNet models pre-trained on the full 1000 classes of ImageNet (generally available from many websites). On the bottom right in blue are accuracy results for what we call “sub-models”, each of which is trained and tested on the same $N_{CL}$, where this number of classes is sampled from the full dataset down to the Few-Class Regime. Findings include the following. (a) Sub-models attain higher upper-bound accuracy than full models. (b) The range of accuracy widens for full models at few-classes, which increases the uncertainty of a practitioner selecting a model for few classes. In contrast, sub-models narrow the range. (c) Full models follow the scaling law (Kaplan et al., 2020)  in the dimension of model size - larger models (darker red) have higher accuracy from many to few classes. (d) Surprisingly, the scaling law is violated for sub-models in the Few-Class Regime (see the zoomed-in subplot) where larger models (darker blue) do not necessarily perform better than smaller ones (lighter blue). From these plots, our key insight is that, instead of using full models, researchers and practitioners in the Few-Class Regime should use sub-models for selection of more efficient models.
@@
We formally define Few-Class Regime as a scenario where the dataset has a limited number of classes. Real-world applications often comprise only a few number of classes (e.g. $N_{CL}<10$ or $10\%$ classes of a dataset). Consequently, Few-Class Arena refers to a benchmark to conduct research experiments to compare models in the Few-Class Regime. This paper focuses on the image classification task, although Few-Class Regime can generalize to object detection (Chen et al., 2019), segmentation (Contributors, 2020) and other visual tasks (described in the Appendix).
@@
  • To the best of our knowledge, we are the first to explore the problems in the Few-Class Regime and develop a benchmark tool Few-Class Arena (FCA) to facilitate scientific research, analysis, and discovery for this range of classes.
@@
  • We incorporate dataset similarity as an inverse difficulty measurement in Few-Class Arena and propose a novel Silhouette-Based Similarity Score named SimSS. By leveraging the visual feature extraction power of CLIP and DINOv2, we show that SimSS is highly correlated with ResNet performance in the Few-Class Regime with high Pearson coefficient scores $\geq 0.88$.
@@ 2 Related Work
Visual Datasets and Benchmarks. To advance deep neural network research, a wealth of large-scale many-class datasets has been developed for benchmarking visual neural networks over a variety of tasks. Typical examples  [^1] A detailed list of many-class datasets used in this paper can be found in the Appendix. include 1000 classes in ImageNet (Deng et al., 2009) for image classification, and 80 object categories in COCO (Lin et al., 2014) for object detection. Previous benchmarks also extend vision to multimodal research such as image-text (Lee et al., 2024; Le et al., 2024; Laurençon et al., 2024; Bitton et al., 2022). While prior works often scale up the number of object categories for general purpose comparison, studies (Fang et al., 2024; Mayo et al., 2023) raise a concern on whether models trained on datasets with such a large number of classes (e.g. ImageNet) can be reliably transferred to real world applications often with far fewer classes. A close work to ours is vision backbone comparison (Goldblum et al., 2024) whose focus is on model architectures. Our perspective differs in a focus on cases with fewer number of classes, which often better aligns with real-world scenarios.
@@
Dataset Difficulty Measurement. Research has shown the existence of inherent dataset difficulty (Mayo et al., 2023) for classification and other analytic tasks. Efficient measurement methods are proposed to characterize dataset difficulty using Silhouette Score (Rousseeuw, 1987), K-means Fréchet inception distance (Dowson & Landau, 1982; Heusel et al., 2017; Lucic et al., 2018), and Probe nets (Scheidegger et al., 2021). Prior studies have proposed image quality metrics using statistical heuristics, including peak signal-to-noise ratio (PSNR) (Hore & Ziou, 2010), structural similarity (SSIM) Index (Wang et al., 2004), and visual information fidelity VIF (Sheikh & Bovik, 2006). A neuroscience-based image difficulty metric (Mayo et al., 2023) is defined as the minimum viewing time related to object solution time (OST) (Kar et al., 2019). Another type of difficulty measure method consists of additional procedures such as c-score (Jiang et al., 2020), prediction depth (Baldock et al., 2021), and adversarial robustness (Goodfellow et al., 2014). Our work aligns with the line of research (Arun, 2012; Trick & Enns, 1998; Wolfe et al., 2010) involving similarity-based difficulty measurements: similar images are harder to distinguish from each other while dissimilar images are easier. Previous studies are mainly in the image retrieval context (Zhang & Lu, 2003; Wang et al., 2014; Tudor Ionescu et al., 2016). Similarity score is used in (Cao et al., 2023) with the limitation that a model serving similarity measurement has to be trained for one dataset. We push beyond this limit by leveraging large vision models that learn general visual features using CLIP (Radford et al., 2021) and DINOv2 (Oquab et al., 2023). The study (Mayo et al., 2023) shows that CLIP generalizes well to both easy and hard images, making it a good candidate for measuring image difficulty. Supported by the evidence that better classifiers can act as better perceptual feature extractors (Kumar et al., 2022), in later sections we show how CLIP and DINOv2 will be used as our similarity base function.
@@
Despite the innovation of difficulty measure algorithms on many-class datasets, little attention has been paid to leveraging these methods in the Few-Class Regime. We show that, as the number of classes decreases, sub-dataset difficulty in the Few-Class Regime plays an increasingly critical role in efficient model selection. To summarize, unlike previous work on many-class benchmarks and difficulty measurements, our work takes few-class and similarity-based dataset difficulty into consideration, and in doing so we believe the work pioneers the development of visual benchmark dedicated to research in the Few-Class Regime.
@@ 3 Few-Class Arena (FCA)
We introduce the Few-Class Arena (FCA) benchmark in this section. In practice, we have integrated FCA into the MMPreTrain framework (Contributors, 2023), implemented in Python and Pytorch [^2] Code is available at https://github.com/bryanbocao/fca, including detailed documentation and long-term plans of maintenance.. Our benchmark usage guidelines are detailed in A.2 of the Appendix.
@@ 3.2 Many-Class Full Dataset Trained Benchmark
Results are summarized in Table 1. We observe (1) models in different datasets (in rows) yield highly variable levels of performance by Top-1 accuracy; (2) no single best model (bold, in columns) exists across all datasets; and (3) model rankings vary across various datasets.
The first two observations are consistent with the findings in (Scheidegger et al., 2021; Fang et al., 2024). For (1), it suggests there exists underlying dataset-specific difficulty. To capture this characteristic, we adopt the reference dataset classification difficulty number (DCN) (Scheidegger et al., 2021) to refer to the empirically highest accuracy achieved in a dataset from a finite number of models shown in Table 1 and Figure 2 (a). For observation (3), we can examine the rankings among the ten datasets of ResNet50 and EfficientNet V2 in Figure 2 (b). ResNet50’s ranking varies dramatically across different datasets, for instance ranking 7th on ImageNet1K and 1st on Quickdraw345. This ranking variability is also observed in other models (see all models in the Appendix). However, a common practice is to benchmark models – even for efficiency – on large datasets, especially ImageNet1K. The varied dataset rankings in our experiments expose the limitations of such a practice, further supporting our new benchmark paradigm, especially in the Few-Class Regime. In later sections, we leverage DCN and image similarity for further analysis.
@@
Table 1: Top-1 accuracy across ten models in ten datasets. Models are trained and tested on full datasets with their original number of classes (e.g. 1K from ImageNet1K) denoted in the last few digits of the abbreviation of the dataset name. The best score is highlighted in bold while the second best is underlined for each dataset. References for all models and datasets are in the Appendix.
@@
In the next subsections, we introduce three new types of benchmarks: (1) Few-Class Full Dataset Trained Benchmark (FC-Full), which benchmarks vision models trained on the full dataset with the original number of classes; (2) Few-Class Subset Trained Benchmark (FC-Sub), which benchmarks vision models trained on subsets of a fewer number of classes than the full dataset, and (3) Few-Class Similarity Benchmark (FC-Sim), which benchmarks image similarity methods and their correlation with model performance.
@@ 3.3 Few-Class Full Dataset Trained Benchmark (FC-Full)
Traditionally, a large number of models are trained and compared on many-class datasets. However, such results cannot be directly transferred to the Few-Class Regime and many real-world scenarios. Therefore, we introduce the Few-Class Full Dataset Trained Benchmark (FC-Full), with the objective of effortlessly conducting large-scale experiments and analyses in the Few-Class Regime.
The procedure of FC-Full consists of two main stages. In the first stage, users select the models and datasets upon which they would like to conduct experiments. They can choose to download pre-trained model weights, which are usually available on popular model hubs (PyTorch Hub (Foundation, 2024), TensorFlow Hub (Inc., 2024), Hugging Face (Face, 2024), MMPreTrain (Contributors, 2023) etc.). In case of no pre-trained weights available from public websites, users can resort to the option of training from scratch. To that end, our tool is designed and implemented to generate bash scripts for easily configurable and modifiable training through the use of configuration files.
In the second stage, users conduct benchmarking in the Few-Class Regime. By specifying the list of classes, Few-Class Arena automatically loads pre-trained weights of the chosen models and evaluates performance of the models on the selected datasets. Note that this process is accomplished through configuration files created by the user’s specifications, thus enabling hundreds of experiments to be launched by a single command. This dramatically reduces the human effort that would otherwise be expended to run these experiments without Few-Class Arena.
@@ 3.5 Few-Class Similarity Benchmark (FC-Sim)
The intrinsic image difficulty of a dataset affects a model’s classification performance (and human) (Geirhos et al., 2017; Rajalingham et al., 2018; Mayo et al., 2023). We show – as is intuitive – that the more similar two images are, the more difficult it is for a vision classifier to make a correct prediction. This suggests that the level of similarity of images in a dataset can be used as a proxy for a dataset difficulty measure. In this section, we first adopt and provide the basic formulation of similarity, the baseline of a similarity metric. Then we propose a Similarity-Based Silhouette Score to capture the characteristic of image similarity in a dataset.
We first adopt the basic similarity formulation from (Cao et al., 2023). Intra-Class Similarity $S_{\alpha}^{(C)}$ is defined as a scalar describing the similarity of images within a class by taking the average of all the distinct class pairs in $C$, while Inter-Class Similarity denotes a scalar describing the similarity among images in two different classes ${C_{1}}$ and $C_{2}$. For a dataset $D$, these are defined as the mean of their similarity scores over all classes, respectively:
$S_{\alpha}^{(D)}=\frac{1}{|L|}\sum_{l\in L}S_{\alpha}^{(C_{l})}=\frac{1}{|L|\times|P^{(C_{l})}|}\sum_{l\in L}\hskip 2.0pt\sum_{i,j\in C_{l};\hskip 2.0pti\neq j}\cos(\textbf{Z}_{i},\textbf{Z}_{j}),$ (1)
$S_{\beta}^{(D)}=\frac{1}{|P^{(D)}|}\sum_{a,b\in L;a\neq b}S_{\beta}^{(C_{a},C_{b})}=\frac{1}{|P^{(D)}|\times|P^{(C_{1},C_{2})}|}\sum_{a,b\in L;\hskip 2.0pta\neq b}\hskip 2.0pt\sum_{i\in C_{1},j\in C_{2}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}),$ (2)
@@
To compensate the aforementioned drawback, we adopt the Silhouette Score (SS) (Rousseeuw, 1987; Shahapure & Nicholas, 2020): $SS(i)=\frac{b(i)-a(i)}{max({a(i),b(i)})}$, where $SS(i)$ is the Silhouette Score of the data point $i$, $a(i)$ is the average dissimilarity between $i$ and other instances in the same class, and $b(i)$ is the average dissimilarity between $i$ and other data points in the closest different class.
Observe that the above Intra-Class Similarity $S_{\alpha}^{(C)}$ already represents the tightness of the class $(C)$, therefore $a(i)$ can be replaced with the inverse of Intra-Class Similarity $a(i)=-S_{\alpha}(i)$. For the second term $b(i)$, we adopt the previously defined Inter-Class Similarity $S_{\beta}^{(C_{1},C_{2})}$ and introduce a new similarity score as Nearest Inter-Class Similarity $S^{\prime(C)}_{\beta}$, which is a scalar describing the similarity among instances between class $C$ and the closest class of each instance in $C$. The dataset-level Nearest Inter-Class Similarity ${S^{\prime}}_{\beta}^{(D)}$ is expressed as:
${S^{\prime}}_{\beta}^{(D)}=\frac{1}{|L|}\sum_{l\in L}{S^{\prime}}_{\beta}^{(C_{l})}=\frac{1}{|L|\times|P^{(C_{l},\hat{C_{l}})}|}\sum_{l\in L}\hskip 2.0pt\sum_{i\in C_{l},j\in\hat{C_{l}}}\cos(\textbf{Z}_{i},\textbf{Z}_{j}),$ (3)
where $\hat{C}$ is the nearest class to instance $i$ ($\hat{C}\neq C$). To summarize, we introduce our novel Similarity-Based Silhouette Score $SimSS$ [^4] The extended derivation is detailed in the Appendix. for dataset $D$: